OpenStudy (anonymous):

Help Please? I just need these checked :) Line 1 (put into slope intercept form)..... (-1, 1) & (0, 3) The answer I had gotten was Y=2x + 3 Line 2 (Slope Intercept form as well)..... (1,5) & (3,3) I had gotten y= -1x + 6. I'm not sure of the second answer but I'm pretty confident that the first one is correct and I would like if someone could just check them.

jimthompson5910 (jim_thompson5910):

here's how you check checking (-1,1) y = 2x + 3 1 = 2(-1) + 3 ... plug in (x,y) = (-1,1) 1 = -2 + 3 1 = 1 .. works --------------- checking (0,3) y = 2x + 3 3 = 2(0) + 3 ... plug in (x,y) = (0,3) 3 = 0 + 3 3 = 3 .. works -------------- both points make the equation true, so the answer is confirmed

jimthompson5910 (jim_thompson5910):

does this make sense?

OpenStudy (anonymous):

Sorta. @jim_thompson5910 I'm gonna check the 2nd one right now

jimthompson5910 (jim_thompson5910):

alright, post your work if possible so i can make sure you're doing it right

OpenStudy (anonymous):

(1, 5) 1 = -1(5) + 6 1 = -5 + 6 1 = 1 . . . 3 = -1(3) + 6 3 = -3 + 6 3 = 3 @jim_thompson5910 Thank you for showing me how to check these! :)

jimthompson5910 (jim_thompson5910):

Thanks for posting your work. Both are correct. Nice job